# Jacobian Descent For Multi-Objective Optimization

### v1

@@ 2.3 Unconflicting projection of gradients
We now define the unconflicting projection of gradients aggregator $\mathcal{A}_{\text{UPGrad}}$, specifically designed to be non-conflicting, linear under scaling and weighted. In essence, it projects each gradient onto the dual cone of the rows of the Jacobian and averages the results, as illustrated in Figure 1(a).
@@
Figure 1: Aggregation of $J=[{\boldsymbol{g}}_{1}\;{\boldsymbol{g}}_{2}]^{\top}\in\mathbb{R}^{2\times 2}$ by four different aggregators. The dual cone of $\{{\boldsymbol{g}}_{1},{\boldsymbol{g}}_{2}\}$ is represented in green.
(a) $\mathcal{A}_{\text{UPGrad}}$ projects ${\boldsymbol{g}}_{1}$ and ${\boldsymbol{g}}_{2}$ onto the dual cone and averages the results.
(b) The mean $\mathcal{A}_{\text{Mean}}(J)=\frac{1}{2}({\boldsymbol{g}}_{1}+{\boldsymbol{g}}_{2})$ conflicts with ${\boldsymbol{g}}_{1}$. $\mathcal{A}_{\text{DualProj}}$ projects this mean onto the dual cone, so it lies on its frontier. $\mathcal{A}_{\text{MGDA}}(J)$ is less aligned with ${\boldsymbol{g}}_{2}$ because of its larger norm.
For any $J\in\mathbb{R}^{m\times n}$ and ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, the projection of ${\boldsymbol{x}}$ onto the dual cone of the rows of $J$ is
$\pi_{J}({\boldsymbol{x}})=\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{y}}\in{\mathbb{R}}^{n}:~{}{\boldsymbol{0}}\preceq J{\boldsymbol{y}}\end{subarray}}\|{\boldsymbol{y}}-{\boldsymbol{x}}\|^{2}.$ (2)
Denoting by ${\boldsymbol{e}}_{i}\in{\mathbb{R}}^{m}$ the $i$th standard basis vector, $J^{\top}{\boldsymbol{e}}_{i}$ is the $i$th row of $J$. $\mathcal{A}_{\text{UPGrad}}$ is defined as
$\mathcal{A}_{\text{UPGrad}}(J)=\frac{1}{m}\sum_{i=1}^{m}\pi_{J}(J^{\top}{\boldsymbol{e}}_{i}).$ (3)
Since the dual cone is convex, any combination of its elements with positive coefficients remains in it. In particular, $\mathcal{A}_{\text{UPGrad}}(J)$ is always in the dual cone of the rows of $J$. $\mathcal{A}_{\text{UPGrad}}$ is thus non-conflicting. Note that if no pair of gradients conflicts, $\mathcal{A}_{\text{UPGrad}}$ simply averages the rows of the Jacobian.
Since $\pi_{J}$ is a projection onto a closed convex cone, if ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ and $0<a\in\mathbb{R}$, then $\pi_{J}(a\cdot{\boldsymbol{x}})=a\cdot\pi_{J}({\boldsymbol{x}})$. By (3), $\mathcal{A}_{\text{UPGrad}}$ is thus linear under scaling.
@@
Let $J\in\mathbb{R}^{m\times n}$. For any ${\boldsymbol{u}}\in{\mathbb{R}}^{m}$, $\pi_{J}(J^{\top}{\boldsymbol{u}})=J^{\top}{\boldsymbol{w}}$ with
${\boldsymbol{w}}\in\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{u}}\preceq{\boldsymbol{v}}\end{subarray}}{\boldsymbol{v}}^{\top}JJ^{\top}{\boldsymbol{v}}.$ (4)
@@
The problem defined in (4) can be solved efficiently using a quadratic programming solver, such as those bundled in qpsolvers [21]. For any $i\in[m]$, let ${\boldsymbol{w}}_{i}$ be given by (4) when substituting ${\boldsymbol{u}}$ with ${\boldsymbol{e}}_{i}$. Then, by Proposition 1,
$\mathcal{A}_{\text{UPGrad}}(J)=J^{\top}\left(\frac{1}{m}\sum_{i=1}^{m}{\boldsymbol{w}}_{i}\right).$ (5)
This provides an efficient implementation of $\mathcal{A}_{\text{UPGrad}}$ and proves that it is weighted. $\mathcal{A}_{\text{UPGrad}}$ can also be easily extended to incorporate a vector of preferences by replacing the averaging in (3) and (5) by a weighted sum with positive weights. This extension remains non-conflicting, linear under scaling and weighted.
@@ 2.4 Convergence
We now prove the convergence of JD with $\mathcal{A}_{\text{UPGrad}}$ when minimizing ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ under standard assumptions. If for a given ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, there exists no ${\boldsymbol{y}}\in{\mathbb{R}}^{n}$ such that ${\boldsymbol{f}}({\boldsymbol{y}})\precneq{\boldsymbol{f}}({\boldsymbol{x}})$, then ${\boldsymbol{x}}$ is said to be Pareto optimal. The set $X^{\ast}\subseteq{\mathbb{R}}^{n}$ of Pareto optimal points is called the Pareto set, and its image ${\boldsymbol{f}}(X^{\ast})$ is called the Pareto front.
Whenever ${\boldsymbol{f}}(\lambda{\boldsymbol{x}}+(1-\lambda){\boldsymbol{y}})\preceq\lambda{\boldsymbol{f}}({\boldsymbol{x}})+(1-\lambda){\boldsymbol{f}}({\boldsymbol{y}})$ holds for any pair of vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$ and any $\lambda\in[0,1]$, ${\boldsymbol{f}}$ is said to be -convex. Moreover, ${\boldsymbol{f}}$ is said to be $\beta$-smooth whenever $\|\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{y}})\|_{\text{F}}\ \varleq\beta\|{\boldsymbol{x}}-{\boldsymbol{y}}\|$ holds for any pair of vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$.
@@
Let ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ be a $\beta$-smooth and -convex function. Suppose that the Pareto front ${\boldsymbol{f}}(X^{\ast})$ is bounded and that for any ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, there is ${\boldsymbol{x}}^{\ast}\in X^{\ast}$ such that ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\preceq{\boldsymbol{f}}({\boldsymbol{x}})$. [^2] This condition is a generalization to the case $m\vargeq 1$ of the condition that there exists a minimizer ${\boldsymbol{x}}^{\ast}\in{\mathbb{R}}^{n}$. Let ${\boldsymbol{x}}_{0}\in{\mathbb{R}}^{n}$, and for all $t\in\mathbb{N}$, ${\boldsymbol{x}}_{t+1}={\boldsymbol{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$, with $\eta=\frac{1}{\beta\sqrt{m}}$. Let ${\boldsymbol{w}}_{t}$ be the weights defining $\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$ as per (5), i.e. $\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t})^{\top}\cdot{\boldsymbol{w}}_{t}$. If ${\boldsymbol{w}}_{t}$ is bounded, then ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})$ for some ${\boldsymbol{x}}^{\ast}\in X^{\ast}$. In other words, ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to the Pareto front.
@@
Experimental considerations suggest that ${\boldsymbol{w}}_{t}$ converges to some ${\boldsymbol{w}}^{\ast}\in{\mathbb{R}}^{m}$ such that both ${\boldsymbol{0}}\prec{\boldsymbol{w}}^{\ast}$ and $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})^{\top}{\boldsymbol{w}}^{\ast}={\boldsymbol{0}}$. This suggests that the boundedness of ${\boldsymbol{w}}_{t}$ could be relaxed or even removed from the set of assumptions of Theorem 1. This would demonstrate that JD with $\mathcal{A}_{\text{UPGrad}}$ and an appropriate step size converges to the Pareto front in the smooth convex case.
@@ 3 Applications
Instance-wise risk minimization
@@
In machine learning, we generally have access to a training set consisting of $m$ examples. The goal of empirical risk minimization (ERM) [18] is simply to minimize the average loss over the whole training set. More generally, instance-wise risk minimization (IWRM) considers the loss associated with each training example as a distinct objective. Formally, if ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ are the parameters of the model and $f_{i}({\boldsymbol{x}})$ is the loss associated to the $i$th example, the respective objective functions of ERM and IWRM are:
(Empirical risk) $\bar{f}({\boldsymbol{x}})$ $=\frac{1}{m}\sum_{i=1}^{m}f_{i}({\boldsymbol{x}})$ (6)
(Instance-wise risk) ${\boldsymbol{f}}({\boldsymbol{x}})$ $=\begin{bmatrix}f_{1}({\boldsymbol{x}})&f_{2}({\boldsymbol{x}})&\cdots&f_{m}({\boldsymbol{x}})\end{bmatrix}^{\top}$ (7)
@@
Naively using GD for ERM is inefficient in most practical cases. A prevalent alternative is to use SGD or one of its variants. Similarly, using JD for IWRM is typically intractable. Indeed, it would require computing a Jacobian matrix with one row per training example at each iteration. In contrast, we can use the Jacobian of a random batch of training example losses. Since it consists of a subset of the rows of the full Jacobian, this approach is a form of stochastic sub-Jacobian descent, as introduced in Section 2.1. Note that IWRM can be extended to cases where each $f_{i}$ is a vector-valued function. The objective would then be the concatenation of the losses of all examples.
Multi-task learning
@@
In multi-task learning, a single model is trained to perform several related tasks simultaneously, leveraging shared representations to improve overall performance [22]. At its core, multi-task learning is a multi-objective optimization problem [10], making it a straightforward application for Jacobian descent. Yet, conflict between tasks is often too limited to justify the overhead of computing all task-specific gradients, i.e. the whole Jacobian [23, 24]. In such cases, a practical approach is to minimize some linear scalarization of the objectives using an SGD-based method. Nevertheless, we believe that a setting with inherent conflict between tasks naturally prescribes Jacobian descent with a non-conflicting aggregator. We analyze several related works applied to multi-task learning in Section 4.
Adversarial training
@@
In adversarial domain adaptation, the feature extractor of a model is trained with two conflicting objectives: The features should be helpful for the main task and should be unable to discriminate the domain of the input [25]. Likewise, in adversarial fairness, the feature extractor is trained to both minimize the predictability of sensitive attributes, such as race or gender, and maximize the performance on the main task [26]. Combining the corresponding gradients with a non-conflicting aggregator could enhance the optimization of such methods. We believe that the training of generative adversarial networks [27] could be similarly formulated as a multi-objective optimization problem. The generator and discriminator could then be jointly optimized with JD.
Momentum-based optimization
@@
In gradient-based single-objective optimization, several methods use some form of gradient momentum to improve their convergence speed [28]. Essentially, their updates consider an exponential moving average of past gradients rather than just the last one. An appealing idea is to modify those algorithms to make them combine the gradient and the momentum with some aggregator, such as $\mathcal{A}_{\text{UPGrad}}$, rather than summing them. This would apply to many popular optimizers, like SGD with Nesterov momentum [29], Adam [19], AdamW [30] and NAdam [31].
Distributed optimization
@@
In a distributed data-parallel setting with multiple machines or multiple GPUs, model updates are computed in parallel. This can be viewed as multi-objective optimization with one objective per data share. Rather than the typical averaging, a specialized aggregator, such as $\mathcal{A}_{\text{UPGrad}}$, could thus combine the model updates. This consideration can even be extended to federated learning, in which multiple entities participate in the training of a common model from their own private data by sharing model updates [32]. In this setting, as security is one of the main challenges, the non-conflicting property of the aggregator could be interesting or even necessary.

### v2

@@ 2.3 Unconflicting projection of gradients
We now define the unconflicting projection of gradients aggregator $\mathcal{A}_{\text{UPGrad}}$, specifically designed to be non-conflicting, linear under scaling, and weighted. In essence, it projects each gradient onto the dual cone of the rows of the Jacobian and averages the results, as illustrated in Figure 1(a).
@@
Figure 1: Aggregation of $J=[{\bm{g}}_{1}\;{\bm{g}}_{2}]^{\top}\in\mathbb{R}^{2\times 2}$ by four different aggregators. The dual cone of $\{{\bm{g}}_{1},{\bm{g}}_{2}\}$ is represented in green.
(a) $\mathcal{A}_{\text{UPGrad}}$ projects ${\bm{g}}_{1}$ and ${\bm{g}}_{2}$ onto the dual cone and averages the results.
(b) The mean $\mathcal{A}_{\text{Mean}}(J)=\frac{1}{2}({\bm{g}}_{1}+{\bm{g}}_{2})$ conflicts with ${\bm{g}}_{1}$. $\mathcal{A}_{\text{DualProj}}$ projects this mean onto the dual cone, so it lies on its boundary. $\mathcal{A}_{\text{MGDA}}(J)$ is almost orthogonal to ${\bm{g}}_{2}$ because of its larger norm.
For any $J\in\mathbb{R}^{m\times n}$ and ${\bm{x}}\in{\mathbb{R}}^{n}$, the projection of ${\bm{x}}$ onto the dual cone of the rows of $J$ is
$\pi_{J}({\bm{x}})=\operatorname*{arg\,min}_{\begin{subarray}{c}{\bm{y}}\in{\mathbb{R}}^{n}:~{}{\bm{0}}\preceq J{\bm{y}}\end{subarray}}\|{\bm{y}}-{\bm{x}}\|^{2}.$ (3)
Denoting by ${\bm{e}_{i}}\in{\mathbb{R}}^{m}$ the $i$th standard basis vector, $J^{\top}{\bm{e}_{i}}$ is the $i$th row of $J$. $\mathcal{A}_{\text{UPGrad}}$ is defined as
$\mathcal{A}_{\text{UPGrad}}(J)=\frac{1}{m}\sum_{i\in[m]}\pi_{J}(J^{\top}{\bm{e}_{i}}).$ (4)
Since the dual cone is convex, it is closed under positive combinations of its elements. For any $J$, $\mathcal{A}_{\text{UPGrad}}(J)$ is thus always in the dual cone of the rows of $J$, so $\mathcal{A}_{\text{UPGrad}}$ is non-conflicting. Note that if no pair of gradients conflicts, $\mathcal{A}_{\text{UPGrad}}$ simply averages the rows of the Jacobian.
Since $\pi_{J}$ is a projection onto a closed convex cone, if ${\bm{x}}\in{\mathbb{R}}^{n}$ and $0<a\in\mathbb{R}$, then $\pi_{J}(a\cdot{\bm{x}})=a\cdot\pi_{J}({\bm{x}})$. By (4), $\mathcal{A}_{\text{UPGrad}}$ is thus linear under scaling.
@@
Let $J\in\mathbb{R}^{m\times n}$. For any ${\bm{u}}\in{\mathbb{R}}^{m}$, $\pi_{J}(J^{\top}{\bm{u}})=J^{\top}{\bm{w}}$ with
${\bm{w}}\in\operatorname*{arg\,min}_{\begin{subarray}{c}{\bm{v}}\in{\mathbb{R}}^{m}:~{}{\bm{u}}\preceq{\bm{v}}\end{subarray}}{\bm{v}}^{\top}JJ^{\top}{\bm{v}}.$ (5)
@@
The problem defined in (5) can be solved efficiently using a quadratic programming solver, such as those bundled in qpsolvers (Caron et al., 2024). For any $i\in[m]$, let ${\bm{w}}_{i}$ be given by (5) when substituting ${\bm{u}}$ with ${\bm{e}_{i}}$. Then, by Proposition 1,
$\mathcal{A}_{\text{UPGrad}}(J)=J^{\top}\left(\frac{1}{m}\sum_{i\in[m]}{\bm{w}}_{i}\right).$ (6)
This provides an efficient implementation of $\mathcal{A}_{\text{UPGrad}}$ and proves that it is weighted. $\mathcal{A}_{\text{UPGrad}}$ can also be easily extended to incorporate a vector of preferences by replacing the average in (4) and (6) by a weighted sum with positive weights. This extension remains non-conflicting, linear under scaling, and weighted.
@@ 2.4 Convergence
We now provide theoretical convergence guarantees of JD with $\mathcal{A}_{\text{UPGrad}}$ when minimizing some ${\bm{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ satisfying standard assumptions. If for a given ${\bm{x}}\in{\mathbb{R}}^{n}$, there exists no ${\bm{y}}\in{\mathbb{R}}^{n}$ satisfying ${\bm{f}}({\bm{y}})\precneq{\bm{f}}({\bm{x}})$, then ${\bm{x}}$ is said to be Pareto optimal. The set $X^{\ast}\subseteq{\mathbb{R}}^{n}$ of Pareto optimal points is called the Pareto set, and its image ${\bm{f}}(X^{\ast})$ is called the Pareto front.
Whenever ${\bm{f}}\big{(}\lambda{\bm{x}}+(1-\lambda){\bm{y}}\big{)}\preceq\lambda{\bm{f}}({\bm{x}})+(1-\lambda){\bm{f}}({\bm{y}})$ holds for any pair of vectors ${\bm{x}},{\bm{y}}\in{\mathbb{R}}^{n}$ and any $\lambda\in[0,1]$, ${\bm{f}}$ is said to be -convex. Moreover, ${\bm{f}}$ is said to be $\beta$-smooth whenever $\big{\|}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})-\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{y}})\big{\|}_{\text{F}}\ \varleq\beta\|{\bm{x}}-{\bm{y}}\|$ holds for any pair of vectors ${\bm{x}},{\bm{y}}\in{\mathbb{R}}^{n}$.
@@
Let ${\bm{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ be a $\beta$-smooth and -convex function. Suppose that the Pareto front ${\bm{f}}(X^{\ast})$ is bounded and that for any ${\bm{x}}\in{\mathbb{R}}^{n}$, there is ${\bm{x}}^{\ast}\in X^{\ast}$ satisfying ${\bm{f}}({\bm{x}}^{\ast})\preceq{\bm{f}}({\bm{x}})$. [^2] This condition is a generalization to the case $m\vargeq 1$ of the existence of a minimizer ${\bm{x}}^{\ast}\in{\mathbb{R}}^{n}$ when $m=1$. Let ${\bm{x}}_{1}\in{\mathbb{R}}^{n}$, and for all $t\vargeq 1$, ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})\big{)}$, with $\eta=\frac{1}{\beta\sqrt{m}}$. Let ${\bm{w}}_{t}$ be the weights defining $\mathcal{A}_{\text{UPGrad}}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})\big{)}$ as per (6), i.e. $\mathcal{A}_{\text{UPGrad}}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})\big{)}=\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})^{\top}\cdot{\bm{w}}_{t}$. If ${\bm{w}}_{t}$ is bounded, then ${\bm{f}}({\bm{x}}_{t})$ converges to ${\bm{f}}({\bm{x}}^{\ast})$ for some ${\bm{x}}^{\ast}\in X^{\ast}$. In other words, ${\bm{f}}({\bm{x}}_{t})$ converges to the Pareto front.
@@
Empirically, ${\bm{w}}_{t}$ appears to converge to some ${\bm{w}}^{\ast}\in{\mathbb{R}}^{m}$ satisfying both ${\bm{0}}\prec{\bm{w}}^{\ast}$ and $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}^{\ast})^{\top}{\bm{w}}^{\ast}={\bm{0}}$. This suggests that the boundedness of ${\bm{w}}_{t}$ could be relaxed or even removed from the set of assumptions of Theorem 1.
Another commonly studied type of convergence for multi-objective optimization is convergence to a stationary point. If for a given ${\bm{x}}\in{\mathbb{R}}^{n}$, there exists ${\bm{0}}\precneq{\bm{w}}$ satisfying $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})^{\top}{\bm{w}}=0$ then ${\bm{x}}$ is said to be Pareto stationary. Even though every Pareto optimal point is Pareto stationary, the converse does not hold, even in the convex case. The function $\begin{bmatrix}x&y\end{bmatrix}^{\top}\mapsto\begin{bmatrix}x^{2}&y^{2}\end{bmatrix}^{\top}$ illustrates this discrepancy. Its Pareto set only contains the origin, but its set of Pareto stationary points is the union of the two axes. Despite being necessary, convergence to a Pareto stationary point is thus not a sufficient condition for optimality and, hence, constitutes a rather weak guarantee. To the best of our knowledge, $\mathcal{A}_{\text{UPGrad}}$ is the first non-conflicting aggregator that provably converges to the Pareto front in the smooth convex case.
@@ 3 Applications
Instance-wise risk minimization.
@@
In machine learning, we generally have access to a training set consisting of $m$ examples. The goal of empirical risk minimization (ERM) (Vapnik, 1995) is simply to minimize the average loss over the whole training set. More generally, instance-wise risk minimization (IWRM) considers the loss associated with each training example as a distinct objective. Formally, if ${\bm{x}}\in{\mathbb{R}}^{n}$ are the parameters of the model and $f_{i}({\bm{x}})$ is the loss associated to the $i$th example, the respective objective functions of ERM and IWRM are:
(Empirical risk) $\bar{f}({\bm{x}})$ $=\frac{1}{m}\sum_{i\in[m]}f_{i}({\bm{x}})$ (7)
(Instance-wise risk) ${\bm{f}}({\bm{x}})$ $=\begin{bmatrix}f_{1}({\bm{x}})&f_{2}({\bm{x}})&\cdots&f_{m}({\bm{x}})\end{bmatrix}^{\top}$ (8)
@@
Naively using GD for ERM is inefficient in most practical cases, so a prevalent alternative is to use SGD or one of its variants. Similarly, using JD for IWRM is typically intractable. Indeed, it would require computing a Jacobian matrix with one row per training example at each iteration. In contrast, we can use the Jacobian of a random batch of training example losses. Since it consists of a subset of the rows of the full Jacobian, this approach is a form of stochastic sub-Jacobian descent, as introduced in Section 2.1. IWRM can also be extended to cases where each $f_{i}$ is a vector-valued function. The objective would then be the concatenation of the losses of all examples.
Multi-task learning.
@@
In multi-task learning, a single model is trained to perform several related tasks simultaneously, leveraging shared representations to improve overall performance (Ruder, 2017). At its core, multi-task learning is a multi-objective optimization problem (Sener & Koltun, 2018), making it a straightforward application for Jacobian descent. Yet, the conflict between tasks is often too limited to justify the overhead of computing all task-specific gradients, i.e. the whole Jacobian (Kurin et al., 2022; Xin et al., 2022). In such cases, a practical approach is to minimize some linear scalarization of the objectives using an SGD-based method. Nevertheless, we believe that a setting with inherent conflict between tasks naturally prescribes Jacobian descent with a non-conflicting aggregator. We analyze several related works applied to multi-task learning in Section 4.
Adversarial training.
@@
In adversarial domain adaptation, the feature extractor of a model is trained with two conflicting objectives: The features should be helpful for the main task and should be unable to discriminate the domain of the input (Ganin et al., 2016). Likewise, in adversarial fairness, the feature extractor is trained to both minimize the predictability of sensitive attributes, such as race or gender, and maximize the performance on the main task (Adel et al., 2019). Combining the corresponding gradients with a non-conflicting aggregator could enhance the optimization of such methods. We believe that the training of generative adversarial networks (Goodfellow et al., 2014) could be similarly formulated as a multi-objective optimization problem. The generator and discriminator could then be jointly optimized with JD.
Momentum-based optimization.
@@
In gradient-based single-objective optimization, several methods use some form of gradient momentum to improve their convergence speed (Polyak, 1964). Essentially, their updates consider an exponential moving average of past gradients rather than just the last one. An appealing idea is to modify those algorithms to make them combine the gradient and the momentum with some aggregator, such as $\mathcal{A}_{\text{UPGrad}}$, instead of summing them. This would apply to many popular optimizers, like SGD with Nesterov momentum (Nesterov, 1983), Adam (Kingma & Ba, 2014), AdamW (Loshchilov & Hutter, 2019) and NAdam (Dozat, 2016).
Distributed optimization.
@@
In a distributed data-parallel setting with multiple machines or multiple GPUs, model updates are computed in parallel. This can be viewed as multi-objective optimization with one objective per data share. Rather than the typical averaging, a specialized aggregator, such as $\mathcal{A}_{\text{UPGrad}}$, could thus combine the model updates. This consideration can even be extended to federated learning, in which multiple entities participate in the training of a common model from their own private data by sharing model updates (Kairouz et al., 2021). In this setting, as security is one of the main challenges, the non-conflicting property of the aggregator could be key.